4.1. You invest R 16 000.00 at an interest rate of 15% per annum. After how many years will your
investment grow to a value of R 40 000.00 if the interest earned is:
4.1.1. simple? (3 marks)
4.1.2. compounded? (5 marks)
4.2. You consider buying a business for R 900 000.
The business is expected to run for 5 years and the following net returns are expected.
YEAR 1 : R 150 000
YEAR 2 : R 200 000
YEAR 3 : R 290 000
YEAR 4 : R 320 000
YEAR 5 : R 220 000
Should you buy the business? A project of this type is expected to return at least 12% per annum. (17 marks)
1
Expert's answer
2019-10-01T10:34:17-0400
4.1.1. 40000 = 16000*(1+0.15*Х)

1+0.15*Х = 2,5

0.15*Х = 1.5

Х = 10 уеars

4.1.2. 40000 = 16000*(1+0.15)x

(1+0.15)x = 2.5

x = 6.56 years


4.2 NPV = 150000/(1+0.12)+ 200000/(1+0.12)2 + 290000/(1+0.12)3 + 320000/(1+0.12)4 + 220000/(1+0.12)5 = 827983.3

You shouldn't  buy business for R 900 000 because 827983.3  less then 900 000